OpenStudy (anonymous):

Completing the Square: x^2+8x-9=0

OpenStudy (anonymous):

(x+4)^2, expand it.

OpenStudy (shubhamsrg):

well theres a logical approach to such questions.. but i'll tell you 1 shortcut.. in ax^2 + bx +c ,, replace x by (y - b/2a) after simplifying,,replace back value of x..

OpenStudy (anonymous):

x^2+8x+16

OpenStudy (anonymous):

Now, how do you make that to be the same as what you started with?

OpenStudy (anonymous):

U understand me? What do you have to do to the expanded equation to get back to the first equation?

OpenStudy (anonymous):

im not sure what I have to do...I am very challenged here.

OpenStudy (anonymous):

Just subtract 25, right?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

16 and 9

OpenStudy (anonymous):

^2+8x+16 x^2+8x-9 Difference 25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So if we take (x+4)^2 and subtract 25 from it we are back to where we started.

OpenStudy (anonymous):

(x+4)^2 -25 = 0 or (x-4)^2 = 25 whci is called completing the square.

OpenStudy (anonymous):

The only tricky bit is coming upo with (x+4)^2 for which you just take half of the number of x's

OpenStudy (anonymous):

couldn't we only do that here because they were perfect squares?

OpenStudy (anonymous):

In theory you can complete the square with any quadratic. In fact, that is how the quadratic formula is derived.

OpenStudy (anonymous):

can we look at another one?

OpenStudy (anonymous):

x^2 + 10x +6 = 0

OpenStudy (anonymous):

so expand x^2+5

OpenStudy (anonymous):

Correct (after you have done a few, you will be able to do it in your head)

OpenStudy (anonymous):

x^2+10x+25

OpenStudy (anonymous):

(x+5)^2

OpenStudy (anonymous):

yeah, I can't do it my head...

OpenStudy (anonymous):

So the difference is?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

So complete the square...

OpenStudy (anonymous):

x^2+10x+19=0

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

or its what the x = x =19

OpenStudy (anonymous):

(x+5)^2 -19 = 0

OpenStudy (anonymous):

what about when the number are not even such as x2-6x+17=0

OpenStudy (anonymous):

(x-3)^2 + 8= 0

OpenStudy (anonymous):

so starting with expanding...x^2+9x+9

OpenStudy (anonymous):

All you are doing is rewriting the original equation as the sum of a square and a constant.

OpenStudy (anonymous):

x^2-6x+9

OpenStudy (anonymous):

yes I messed that one up

OpenStudy (anonymous):

Last one, x^2+ 9x -40 =0

OpenStudy (anonymous):

my teacher put the -8 on a square root? W an answer of x=3+- i square root 8...??

OpenStudy (anonymous):

sqrt(ab) = sqrt(a)*sqrt(b) so Sqrt -8 = Sqrt(-1)*Sqrt(8) = i*8 (or 8i)

OpenStudy (anonymous):

are these just logistics of how to write the answer?

OpenStudy (anonymous):

I wrote that last wrong sqrt(ab) = sqrt(a)*sqrt(b) so Sqrt -8 = Sqrt(-1)*Sqrt(8) = i*sqrt8 (or sqrt(8)i) Logistics?

OpenStudy (anonymous):

ok so you would write it 3 +- i sq rt 8 ?

OpenStudy (anonymous):

That's OK but usually you reduce the sqrt 8 to 2sqrt2

OpenStudy (anonymous):

Teacher will probably dock you a mark if you don't reduce it.

OpenStudy (anonymous):

ok so the last example you gave...not sure what to do now if the number isn't even...

OpenStudy (anonymous):

Exactly the same as before, just take the half of it

OpenStudy (anonymous):

(x +9/2)^2

OpenStudy (anonymous):

And if it was x^2 + bx +c you still do it just the same (x+b/2)^2

OpenStudy (anonymous):

Anyway, I leave you to practice...:-)

OpenStudy (anonymous):

thank you very much!